OpenStudy (anonymous):

ALG 2... PLEASE HELP! ILL GIVE A MEDAL! ^.^ What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 36 y = x - 6 2 and 0 6 and 2 6 and -2 6 and 0

OpenStudy (mathstudent55):

Since the second equation is already solved for y, use the substitution method.

OpenStudy (anonymous):

welp, i know its not 6 and 2...

OpenStudy (mathstudent55):

@kmeis002 In this problem, you can't just plug in points to see if they work, because of the wording of the question.

OpenStudy (anonymous):

To solve that original syatem of equations: First note that we take y = x - 6 and substiutute that into y in the first equation, So the FIRST equation becomes x^2 + (x-6)^2 = 36 Simplify the above, you get x^2 + x^2 - 12x + 36 = 36 2x^2 - 12 x = 0 2x(x - 6) = 0 x = 0 or x = 6 If x = 0, y = -6; if x = 6, y = 0 Final solutions: x = 0 and y = -6 AND x = 6 and y= 0

OpenStudy (kmeis002):

Ah I see, thank you @mathstudent55, I had misread.

OpenStudy (mathstudent55):

\(x^2 + y^2 = 36\) \(y = x - 6\) \(x^2 + \color{red}{y}^2 = 36\) \(y = \color{red}{x - 6}\) \(x^2 + \color{red}{(x - 6)}^2 = 36\) That is the substitution step.Now solve the bottom equation for x.

OpenStudy (anonymous):

i still cant get it.. .-.

OpenStudy (mathstudent55):

Now we solve that quadratic equation. First, we square the binomial x - 6: \(x^2 + (x - 6)^2 = 36\) \(x^2 + x^2 - 12x + 36 = 36\) \(2x^2 - 12x + 36 = 36\) \(2x^2 - 12x = 0\) \(x^2 - 6x = 0\) \(x(x - 6) = 0\) x = 0 or x - 6 = 0 x = 0 or x = 6

OpenStudy (anonymous):

ohh, okay thank you!!

OpenStudy (mathstudent55):

You're welcome.